OpenStudy (anonymous):

Calc Question, Will Medal! Find y' if y=(x^e)(e^x) I understand the first portion of the problem (product rule) but having trouble with other parts.

OpenStudy (anonymous):

\[y=x^{e}e^{x}\]

zepdrix (zepdrix):

Oo that's a cool problem :)

OpenStudy (anonymous):

Hahaha

zepdrix (zepdrix):

Start with product rule, ya?\[\large\rm y'=(x^e)'e^x+x^e(e^x)'\]

zepdrix (zepdrix):

Derivative of e^x?

OpenStudy (anonymous):

The derivative of e^x would be e^x

zepdrix (zepdrix):

\[\large\rm y'=(x^e)'e^x+x^e(e^x)\]Ok cool.

OpenStudy (anonymous):

I am franticly looking through my notes....

OpenStudy (anonymous):

Sorry :(

zepdrix (zepdrix):

It's simpler than you think :) \(\large\rm \frac{d}{dx}x^n=nx^{n-1}\)

OpenStudy (anonymous):

Ohhhhh!!! So that means I can simply do: \[y' = (ex ^{e-1})(e ^{x}) + (e ^{x})(x ^{e})\]

zepdrix (zepdrix):

Yes, good job. From that point, you could do some factoring, not completely necessary though.

OpenStudy (anonymous):

Ok, would I next have to get the natural log of parts of the equation?

OpenStudy (anonymous):

I take that back... that is pretty much the answer isn't it?

zepdrix (zepdrix):

Mmmmm, no. No logs required. But as I said, you could factor if you like,\[\large\rm y'=ex^{e-1}e^x+x^ee^x\]\[\large\rm y'=x^{e-1}e^x(e+x)\]

zepdrix (zepdrix):

Yes, you pretty much have your answer :))

OpenStudy (anonymous):

Haha thank you so much for you time. I appreciate it!

zepdrix (zepdrix):

np